OpenStudy (anonymous):

A hemispherical dome with radius 50 feet will be given a coat of paint 0.01 inch thick. The contractor for the job wants to estimate the number of gallons of paint that will be needed. Use differentials to obtain an estimate (there are 231 cubic inches in a gallon). HINT:Approximate the change in the volume of the hemisphere corresponding to an increase of 0.01 inch in the radius.

OpenStudy (anonymous):

volume of a 1/2 sphere = (2/3)pi r^3 dV = 2pi r^2 dr r = 50ft = 600in dr=0.01in dV = 2 pi (600^2)(0.01) = (some value) cubic inches (i dont have a calculator) no. of gallons = (some value)/(231 cubic in per gallon)

OpenStudy (anonymous):

yep i ending up getting 98 gallons